If the ratio of 3s to 4t is 1/2, what is the ratio of 5t to s?
Answer by John10(297)

If the ratio of 3s to 4t is 1/2, what is the ratio of 5t to s?
--------------------------
First part: 3s/ 4t = 1/2
Then 6s = 4t
s = 4t/6
s = 2t/3
ratio of 5t to s or 5t/s = ?
You can replace s of 5t/s by 2t/3 to find the ratio.
Best of luck! John